[Brahms-l] Shift report 20050310 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Mar 10 2005 - 16:05:31 EST
Shifter leader was JHL
Shifters were 


Beam for 2.5 hours
  -10 am (ZDC at the beginning of store: 1100/sec)
  3:30pm- (ZDC at the beginning of store: 200/sec)

Data run: MRS@90 4(B)KG FS@8 1/4A

T3,T4 have been fixed by Bob. (See e-log)
See elog for some numbers for Spec trigger efficiencies.




_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 10 16:25:57 2005

This archive was generated by hypermail 2.1.8 : Thu Mar 10 2005 - 16:26:06 EST